OpenStudy (anonymous):

8√45 - 3√80 simplify

OpenStudy (solomonzelman):

45=9 * 5 80= 16 * 5 you know that 9 and 16 are perfect squares.

OpenStudy (solomonzelman):

45=9 * 5 80= 16 * 5 you know that 9 and 16 are perfect squares.

OpenStudy (anonymous):

so it would be 9√5 - 16√5? then simplify?

OpenStudy (solomonzelman):

\(\Large\color{blueviolet }{ \sf \sqrt{45} =\sqrt{9 \times 5 }=\sqrt{9 } \times\sqrt{ 5 }=3\sqrt{ 5 }}\) \(\Large\color{blue }{ \sf \sqrt{80} =\sqrt{16\times 5 }=\sqrt{16 } \times\sqrt{ 5 }=4\sqrt{ 5 }}\) \(\Large\color{green}{ \sf 8\sqrt{45} -3\sqrt{80} =8~(3\sqrt{ 5 })-3~(4\sqrt{ 5 })}\) finish it off please.

OpenStudy (anonymous):

i dont understand how you got the 3rd part

OpenStudy (solomonzelman):

the third part, is the begging on how to simplify based on what we know from blue and purple.

OpenStudy (anonymous):

i dont understand what the final answer would be

OpenStudy (solomonzelman):

\(\Large\color{red}{ \sf 8~(3\sqrt{ 5 })-3~(4\sqrt{ 5 })=}\) \(\Large\color{red}{ \sf 24\sqrt{ 5 }-12\sqrt{ 5 }= ?}\)

OpenStudy (anonymous):

12√5

OpenStudy (solomonzelman):

Yes, it would be 12√5 Also, I want to make your life easier (you probably don't know this) 1) click ALT and hold it 2) click 2 5 1 3) release the ALT try it.

OpenStudy (anonymous):

i just copy paste it lol

OpenStudy (solomonzelman):

but click the 2 5 1 on your keyboard on the right, NOT the ones below F1 F2 ..

OpenStudy (anonymous):

i got it but i usually just copy paste it

OpenStudy (anonymous):

thanks man you've helped me alot

OpenStudy (solomonzelman):

Well, it's certainly easier :) you welcome.